# rank-33 generalized Clifford manifold and its twistor space

### 10 Mar 2026 v1

@@ Abstract.
We introduce the notion of a rank-3 generalized Clifford manifold, defined by a triple of generalized complex structures satisfying Clifford-type relations. We show that every such structure canonically induces a generalized hypercomplex structure. We further describe a natural Spin(3)-action by Clifford rotations, which produces an $S^{2}\times S^{2}$-family of generalized complex structures. The corresponding twistor space is then constructed, and we prove that the induced almost generalized complex structure is integrable. In contrast to the standard pure-spinor approach, the integrability of the twistor-space structure is established entirely in terms of the generalized Nijenhuis tensor.
@@ 1. Introduction
Let $\left(\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}\right)$ be an alomost generalized Clifford structure over $\mathbb{T}M$, then
@@
where $\mathcal{J}_{i}$ is the induced almost generalized complex structure defined by (1.1),
We further show that rank-$3$ generalized Clifford structures admit natural $Spin(3)$-transformations induced by orthogonal rotations. These transformations act as rank-$3$ Clifford rotations and generate an $S^{2}\times S^{2}$-family of generalized complex structures, extending the classical twistor sphere construction. Via stereographic projection, each Riemann sphere is identified with $S^{2}\subset\mathbb{R}^{3}$. The associated rotation matrices $T\left(\zeta_{1}\right)$ and $S\left(\zeta_{2}\right)\in SO(3)$ encode the Clifford rotations determined by the parameters $\zeta_{1}$ and $\zeta_{2}$, which are the stereographic complex coordinates on the two factors of $S^{2}\times S^{2}$. The explicit formulas for these matrices are provided in Section $3$. So we establish a twistor transformation for generalized Clifford manifolds, showing that Clifford rotations generate a natural $S^{2}$-family of generalized complex structures over $TM\oplus T^{*}M$.
@@
is a $Spin(3)$-transformation, which is a rank-$3$ Clifford rotation. And the triple
@@
In Sect. 2, we investigate several structural properties of generalized complex structures. These results play a crucial role in proving the integrability of the generalized complex structure on the twistor space. In Sect. 3, we present a detailed definition of generalized Clifford structures and provide several classical examples. In Sect. 4, we further introduce the $Spin(3)$-transformations associated with rank-$3$ generalized Clifford structures and construct the twistor space of a generalized Clifford manifold.
@@ 2. Preliminaries
(2.1) $X+\xi,Y+\eta]_{D}=[X,Y]+L_{X}\eta-\iota_{Y}{\rm d}\xi,$
which is Jacobian but non-skew. The twisted version of a bracket $X+\xi,Y+\eta]_{D}$ over $\mathbb{T}M$ with a $3$-form $H$ is defined by
@@
$2\mathcal{N}(\mathcal{I},\mathcal{J})\left(A,B\right)=$ $\mathcal{N}_{\mathcal{I}\mathcal{J}}\left(\mathcal{J}A,\mathcal{I}B\right)+\mathcal{I}\mathcal{J}\left(\mathcal{N}_{\mathcal{I}}\left(A,B\right)-\mathcal{N}_{\mathcal{J}}\left(A,B\right)\right),$
@@
The generalized matric $\mathcal{G}$, viewed as a generalized almost real structure, is in general not integrable.
@@ 3. Generalized Clifford manifold
The induced almost generalized complex structure $\mathcal{J}_{i}=\frac{1}{2}\epsilon_{ijk}\mathcal{I}_{j}\mathcal{I}_{k}$ in (1.1) is integrable. Then $\left(\mathcal{J}_{1},\mathcal{J}_{2},\mathcal{J}_{3}\right)$ forms a generalized hypercomplex structure and the generalized Clifford manifold must be generalized hypercomplex.
@@
(3.11) $\begin{array}[]{ll}\mathcal{I}_{i}\mathcal{I}_{j}=-\delta_{ij}1_{2d}+\epsilon_{ijk}\mathcal{I}_{k},&\tilde{\mathcal{I}}_{i}{\tilde{\mathcal{I}}_{j}}=-\delta_{ij}1_{2d}+\epsilon_{ijk}\mathcal{I}_{k},\\ \mathcal{I}_{i}\tilde{\mathcal{I}}_{j}=-\delta_{ij}{G}+\epsilon_{ijk}\tilde{\mathcal{I}}_{k},&\tilde{\mathcal{I}}_{i}\mathcal{I}_{j}=-\delta_{ij}{G}+\epsilon_{ijk}\tilde{\mathcal{I}}_{k},\end{array}$
@@
The generalized Clifford manifold $\left(M,\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}\right)$ is generalized hyperkähler, if $\mathcal{G}=-\mathcal{I}_{1}\mathcal{I}_{2}\mathcal{I}_{3}$ is a generalized metric.
@@
4. the twistor space of generalized Clifford manifold
@@
(4.11) $\mathcal{K}_{i}(\zeta_{1},\zeta_{2}):=\frac{1}{2}(\hat{\mathcal{I}}^{+}_{i}+\hat{\mathcal{I}}^{-}_{i})=\sum_{l=1}^{3}\left(t_{il}\mathcal{I}_{l}^{+}+s_{il}\mathcal{I}_{l}^{-}\right),\ i=1,2,3.$
@@
By the bilinearity of Nijenhuis vector, we have
@@
where $V_{\xi_{1}}^{0,1}$ depends solely on $\left(\zeta_{1},\bar{\zeta}_{1}\right)$ and is proportional to $\mathrm{d}\bar{\zeta}_{1}$, whereas $V_{\zeta_{2}}^{0,1}$ depends solely on $\left(\zeta_{2},\bar{\zeta}_{2}\right)$ and is proportional to $\mathrm{d}\bar{\zeta}_{2}$. This follows from the definition $V^{0,1}=\tfrac{1}{2}\Omega^{0,1}$ and from (4.30) because $\vec{c}$ depends only on $\zeta_{1}$ and $\vec{d}$ only on $\zeta_{2}$. Hence $V^{0,1}_{\zeta_{1}}=A(\zeta_{1},\bar{\zeta}_{1})\,{\rm d}\bar{\zeta}_{1}$ and $V^{0,1}_{\zeta_{2}}=B(\zeta_{2},\bar{\zeta}_{2})\,{\rm d}\bar{\zeta}_{2}$, where

### 15 Mar 2026 v2

@@ Abstract.
We introduce the notion of a rank-3 generalized Clifford manifold, defined by a triple of generalized complex structures satisfying Clifford-type relations. We show that every such structure canonically induces a generalized hypercomplex structure. We further describe a natural Spin(3)-action by Clifford rotations, which produces an $S^{2}\times S^{2}$-family of generalized complex structures. The corresponding twistor space is then constructed, and we prove that the induced almost generalized complex structure is integrable. In contrast to the standard pure-spinor approach, the integrability of the twistor-space structure is established entirely in terms of the generalized Nijenhuis tensor. We further prove that this Clifford-to-twistor construction is compatible with T-duality, in the sense that T-duality preserves the rank-3 Clifford triple, the induced structures, and the associated Spin(3)-rotated family.
@@ 1. Introduction
Let $\left(\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}\right)$ be an almost generalized Clifford structure on $\mathbb{T}M:=TM\oplus T^{*}M$, then
@@
where $\mathcal{J}_{i}$ is the induced almost generalized complex structure defined by (1.1).
We further show that rank-$3$ generalized Clifford structures admit natural $Spin(3)$-transformations induced by orthogonal rotations. These transformations act as rank-$3$ Clifford rotations and generate an $S^{2}\times S^{2}$-family of generalized complex structures, extending the classical twistor sphere construction. Via stereographic projection, each Riemann sphere is identified with $S^{2}\subset\mathbb{R}^{3}$. The associated rotation matrices $T\left(\zeta_{1}\right)$ and $S\left(\zeta_{2}\right)\in SO(3)$ encode the Clifford rotations determined by the parameters $\zeta_{1}$ and $\zeta_{2}$, which are the stereographic complex coordinates on the two factors of $S^{2}\times S^{2}$. The explicit formulas for these matrices are provided in Section 4. So we establish a twistor transformation for generalized Clifford manifolds, showing that Clifford rotations generate a natural $S^{2}\times S^{2}$-family of generalized complex structures over $TM\oplus T^{*}M$.
@@
defines a rank-3 Clifford rotation. Moreover, the triple
@@
In Sect. 2, we investigate several structural properties of generalized complex structures in terms of the generalized Nijenhuis tensor. These results play a crucial role in proving the integrability of the generalized complex structure on the twistor space. In Sect. 3, we introduce generalized Clifford structures and present several basic examples. In Sect. 4, we construct the Spin(3)-transformations associated with rank-3 generalized Clifford structures and establish the generalized complex structure on the twistor space. In Sect. 5, we further study the behavior of rank-3 twisted generalized Clifford structures under T-duality, showing that T-duality preserves not only the original Clifford triple, but also the induced structures and the associated Spin(3)-rotated family.
@@ 2. Preliminaries
(2.1) $X+\xi,Y+\eta]_{D}=[X,Y]+L_{X}\eta-\iota_{Y}{\rm d}\xi.$
The Dorfman bracket satisfies the Jacobi identity but is not skew-symmetric. The twisted version of a bracket $X+\xi,Y+\eta]_{D}$ over $\mathbb{T}M$ with a $3$-form $H$ is defined by
@@ 2.1. Generalized complex structure
$2\mathcal{N}(\mathcal{I},\mathcal{J})\left(A,B\right)=$ $-\mathcal{N}_{\mathcal{I}\mathcal{J}}\left(\mathcal{J}A,\mathcal{I}B\right)+\mathcal{I}\mathcal{J}\left(\mathcal{N}_{\mathcal{I}}\left(A,B\right)-\mathcal{N}_{\mathcal{J}}\left(A,B\right)\right),$
@@ 2.2. Generalized real structure
The generalized metric $\mathcal{G}$, viewed as a generalized almost real structure, is in general not integrable.
@@ 3. Generalized Clifford manifold
The induced almost generalized complex structure $\mathcal{J}_{i}=\frac{1}{2}\epsilon_{ijk}\mathcal{I}_{j}\mathcal{I}_{k}$ in (1.1) is integrable. Then the generalized Clifford manifold canonically determines a generalized hypercomplex structure $\left(\mathcal{J}_{1},\mathcal{J}_{2},\mathcal{J}_{3}\right)$.
@@
(3.11) $\begin{array}[]{ll}\mathcal{I}_{i}\mathcal{I}_{j}=-\delta_{ij}{\rm Id}+\epsilon_{ijk}\mathcal{I}_{k},&\tilde{\mathcal{I}}_{i}{\tilde{\mathcal{I}}_{j}}=-\delta_{ij}{\rm Id}+\epsilon_{ijk}\mathcal{I}_{k},\\ \mathcal{I}_{i}\tilde{\mathcal{I}}_{j}=-\delta_{ij}{G}+\epsilon_{ijk}\tilde{\mathcal{I}}_{k},&\tilde{\mathcal{I}}_{i}\mathcal{I}_{j}=-\delta_{ij}{G}+\epsilon_{ijk}\tilde{\mathcal{I}}_{k},\end{array}$
@@
The generalized Clifford manifold $\left(M,\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}\right)$ induces a generalized hyperkähler structure when $\mathcal{G}=-\mathcal{I}_{1}\mathcal{I}_{2}\mathcal{I}_{3}$ is a generalized metric.
@@
4. The twistor space of a generalized Clifford manifold
@@
(4.11) $\mathcal{K}_{i}(\zeta_{1},\zeta_{2}):=\hat{\mathcal{I}}^{+}_{i}+\hat{\mathcal{I}}^{-}_{i}=\sum_{l=1}^{3}\left(t_{il}\mathcal{I}_{l}^{+}+s_{il}\mathcal{I}_{l}^{-}\right),\ i=1,2,3.$
@@
By the bilinearity of Nijenhuis tensor, we have
@@
where $V_{\zeta_{1}}^{0,1}$ depends solely on $\left(\zeta_{1},\bar{\zeta}_{1}\right)$ and is proportional to $\mathrm{d}\bar{\zeta}_{1}$, whereas $V_{\zeta_{2}}^{0,1}$ depends solely on $\left(\zeta_{2},\bar{\zeta}_{2}\right)$ and is proportional to $\mathrm{d}\bar{\zeta}_{2}$. This follows from the definition $V^{0,1}=\tfrac{1}{2}\Omega^{0,1}$ and from (4.30) because $\vec{c}$ depends only on $\zeta_{1}$ and $\vec{d}$ only on $\zeta_{2}$. Hence $V^{0,1}_{\zeta_{1}}=A(\zeta_{1},\bar{\zeta}_{1})\,{\rm d}\bar{\zeta}_{1}$ and $V^{0,1}_{\zeta_{2}}=B(\zeta_{2},\bar{\zeta}_{2})\,{\rm d}\bar{\zeta}_{2}$, where
@@
5. T-duality and rank-3 generalized Clifford structures
The formulation of T-duality used here follows the generalized-geometric approach of Cavalcanti and Gualtieri [5], building on the topological T-duality framework for circle and torus bundles with $H$-flux developed in [2, 3]. In this section, we show that rank-3 twisted generalized Clifford structures are stable under T-duality. More precisely, we prove that T-duality preserves not only the original Clifford triple $(\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3})$, but also the induced structures $(\mathcal{J}_{1},\mathcal{J}_{2},\mathcal{J}_{3},\mathcal{G})$ and the associated Spin(3)-rotated family $(\mathcal{K}_{1}(\zeta_{1},\zeta_{2}),\mathcal{K}_{2}(\zeta_{1},\zeta_{2}),\mathcal{K}_{3}(\zeta_{1},\zeta_{2}))$. This shows that the Clifford-to-twistor construction developed in the previous sections is compatible with the natural symmetry of generalized geometry given by T-duality.
[…]
$\Phi:\mathbb{T}M=TM\oplus T^{*}M\longrightarrow\mathbb{T}\widetilde{M}=T\widetilde{M}\oplus T^{*}\widetilde{M}$
such that
(5.1) $\langle\Phi(A),\Phi(B)\rangle=\langle A,B\rangle,$
and
(5.2) $\Phi\bigl([A,B]_{H}\bigr)=[\Phi(A),\Phi(B)]_{\widetilde{H}},\qquad A,B\in\Gamma(\mathbb{T}M).$
Here $\cdot,\cdot]_{H}$ and $\cdot,\cdot]_{\widetilde{H}}$ denote the $H$-twisted and $\widetilde{H}-$twisted Dorfman brackets, respectively.
Condition (5.2) implies that any generalized-geometric structure defined in terms of the natural pairing and the twisted Dorfman bracket can be transported through $\Phi$. In particular, for any bundle endomorphism $\mathcal{I}\in\Gamma(End(\mathbb{T}M))$, we define its T-dual transform by
$).\widetilde{\mathcal{I}}=\Phi\mathcal{I}\Phi^{-1}\in\Gamma(End(\mathbb{T}\widetilde{M}$)).
Lemma 5.1.
Let $\mathcal{I},\mathcal{J}\in\Gamma(End(\mathbb{T}M))$. Then the mixed generalized Nijenhuis tensor is preserved by T-duality in the sense that
$\mathcal{N}_{\widetilde{H}}(\widetilde{\mathcal{I}},\widetilde{\mathcal{J}})(\Phi A,\Phi B)=\Phi\Bigl(\mathcal{N}_{H}(\mathcal{I},\mathcal{J})(A,B)\Bigr),\qquad A,B\in\Gamma(\mathbb{T}M).$
In particular,
$\mathcal{N}_{H}(\mathcal{I},\mathcal{J})=0\quad\Longrightarrow\quad\mathcal{N}_{\widetilde{H}}(\widetilde{\mathcal{I}},\widetilde{\mathcal{J}})=0.$
Proof.
This follows directly from the definition of the generalized Nijenhuis concomitant and the fact that $\Phi$ intertwines the twisted Dorfman brackets:
$\Phi([A,B]_{H})=[\Phi(A),\Phi(B)]_{\widetilde{H}}$.
Indeed, every term in the definition of $\mathcal{N}(\mathcal{I},\mathcal{J})$ is functorial under conjugation by $\Phi$, and therefore the whole expression is transported to the corresponding one on $\mathbb{T}\widetilde{M}$. ∎
We now show that T-duality preserves the class of rank-3 twisted generalized Clifford manifolds.
Proposition 5.1.
[…]
be an orthogonal Courant algebroid isomorphism. Assume that $(\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3})$ is an $H$-twisted rank-3 generalized Clifford structure on $M$. Define
$\widetilde{\mathcal{I}}_{a}:=\Phi\mathcal{I}_{a}\Phi^{-1},\qquad a=1,2,3.$
Then $(\widetilde{I}_{1},\widetilde{I}_{2},\widetilde{I}_{3})$ is a $\widetilde{H}-$twisted rank-3 generalized Clifford structure on $\widetilde{M}$.
Proof.
Since each $\mathcal{I}_{a}$ is orthogonal and satisfies $\mathcal{I}_{a}^{2}=-{\rm Id}$, and since $\Phi$ is orthogonal, each $\widetilde{I}_{a}$ is again an orthogonal almost generalized complex structure on $\mathbb{T}\widetilde{M}$:
$\widetilde{\mathcal{I}}_{a}^{2}=\Phi\mathcal{I}_{a}^{2}\Phi^{-1}=-{\rm Id}.$
Moreover, the Clifford relations are preserved:
$\widetilde{\mathcal{I}}_{a}\widetilde{\mathcal{I}}_{b}+\widetilde{\mathcal{I}}_{b}\widetilde{\mathcal{I}}_{a}=\Phi(\mathcal{I}_{a}\mathcal{I}_{b}+\mathcal{I}_{b}\mathcal{I}_{a})\Phi^{-1}=-2\delta_{ab}{\rm Id}.$
Finally, since each $I_{a}$ is integrable with respect to the $H$-twisted Dorfman bracket, Lemma 5.1 gives
$\mathcal{N}_{\widetilde{H}}(\widetilde{I}_{a},\widetilde{I}_{a})=0,\qquad a=1,2,3.$
Hence $(\widetilde{I}_{1},\widetilde{I}_{2},\widetilde{I}_{3})$ defines a $\widetilde{H}-$twisted rank-3 generalized Clifford structure on $\widetilde{M}$. ∎
We now show that these induced structures (1.1) are also compatible with T-duality.
Proposition 5.2.
With the notation (1.1), define
(5.3) $\widetilde{\mathcal{J}}_{i}:=\frac{1}{2}\varepsilon_{ijk}\widetilde{\mathcal{I}}_{j}\widetilde{\mathcal{I}}_{k},\qquad\widetilde{\mathcal{G}}:=-\widetilde{\mathcal{I}}_{1}\widetilde{\mathcal{I}}_{2}\widetilde{\mathcal{I}}_{3}.$
Then
$\widetilde{\mathcal{J}}_{i}=\Phi\mathcal{J}_{i}\Phi^{-1},\qquad\widetilde{\mathcal{G}}=\Phi\mathcal{G}\Phi^{-1}.$
Proof.
Using the definition of $\mathcal{J}_{i}$ in (1.1) and the multiplicativity of conjugation, we obtain
$\widetilde{\mathcal{J}}_{i}=\frac{1}{2}\varepsilon_{ijk}\widetilde{\mathcal{I}}_{j}\widetilde{\mathcal{I}}_{k}=\frac{1}{2}\varepsilon_{ijk}(\Phi\mathcal{I}_{j}\Phi^{-1})(\Phi\mathcal{I}_{k}\Phi^{-1})=\Phi\Bigl(\frac{1}{2}\varepsilon_{ijk}\mathcal{I}_{j}\mathcal{I}_{k}\Bigr)\Phi^{-1}=\Phi\mathcal{J}_{i}\Phi^{-1}.$
Similarly,
$\widetilde{\mathcal{G}}=-\widetilde{\mathcal{I}}_{1}\widetilde{\mathcal{I}}_{2}\widetilde{\mathcal{I}}_{3}=-\Phi\mathcal{I}_{1}\mathcal{I}_{2}\mathcal{I}_{3}\Phi^{-1}=\Phi\mathcal{G}\Phi^{-1}.$
As an immediate consequence, all algebraic identities in (1.2) are preserved by T-duality. Define similarly
[…]
and let
$\bigl(\widetilde{\mathcal{K}}_{1}(\zeta_{1},\zeta_{2}),\widetilde{\mathcal{K}}_{2}(\zeta_{1},\zeta_{2}),\widetilde{\mathcal{K}}_{3}(\zeta_{1},\zeta_{2})\bigr)$
be the corresponding Clifford-rotated family on $\widetilde{M}$.
Proposition 5.3.
For every $(\zeta_{1},\zeta_{2})\in\mathbb{CP}^{1}\times\mathbb{CP}^{1}$, one has
$\widetilde{\mathcal{K}}_{i}(\zeta_{1},\zeta_{2})=\Phi{\mathcal{K}}_{i}(\zeta_{1},\zeta_{2})\Phi^{-1},\qquad i=1,2,3.$
Equivalently, T-duality commutes with the rank-3 Clifford rotations.
Proof.
By Proposition 5.2, we have
$\widetilde{\mathcal{I}}_{i}^{\pm}=\frac{1}{2}(\widetilde{\mathcal{J}}_{i}\pm\widetilde{\mathcal{I}}_{i})=\frac{1}{2}\Phi({\mathcal{J}}_{i}\pm{\mathcal{I}}_{i})\Phi^{-1}=\Phi{\mathcal{I}}_{i}^{\pm}\Phi^{-1}.$
Hence, using the definition of the rotated structures in Section 4,
$\widehat{\widetilde{\mathcal{I}}}_{i}^{+}=\sum_{l=1}^{3}t_{il}\widetilde{\mathcal{I}}_{l}^{+}=\sum_{l=1}^{3}t_{il}\Phi I_{l}^{+}\Phi^{-1}=\Phi\Bigl(\sum_{l=1}^{3}t_{il}{\mathcal{I}}_{l}^{+}\Bigr)\Phi^{-1}=\Phi\widehat{\mathcal{I}}_{i}^{\,+}\Phi^{-1},$
and similarly
$\widehat{\widetilde{\mathcal{I}}}_{i}^{-}=\Phi\widehat{\mathcal{I}}_{i}^{\,-}\Phi^{-1},$
Therefore,
$\widetilde{\mathcal{K}}_{i}=\frac{1}{2}\bigl(\widetilde{\widehat{I}}_{i}^{\,+}+\widetilde{\widehat{I}}_{i}^{\,-}\bigr)=\Phi\Bigl(\frac{1}{2}(\widehat{{\mathcal{I}}}_{i}^{+}+\widehat{{\mathcal{I}}}_{i}^{-})\Bigr)\Phi^{-1}=\Phi\mathcal{K}_{i}\Phi^{-1}.$
This proves the claim. ∎
Combining Theorem 5.1 with Proposition 5.3, we obtain:
Corollary 5.1.
The T-dual of the Spin(3)-rotated family associated with a rank-3 twisted generalized Clifford structure is exactly the Spin(3)-rotated family associated with the T-dual Clifford triple. In particular, the construction of the $S^{2}\times S^{2}$-family of generalized complex structures is natural with respect to T-duality.
Remark 5.1.
Proposition 5.1 and Proposition 5.3 show that the $M$-component of the twistor generalized complex structure constructed in Theorem 1.3 is compatible with T-duality. It is therefore natural to expect that, after pulling back the twist to $M\times S^{2}\times S^{2}$, the corresponding generalized complex structure on the twistor space admits a T-dual interpretation as well. We leave this total-space formulation to future work.